Two firms, Monolyte and NuPlant, are emitting toxic smoke into the local town. Monolyte emits 5 tons, each of which would cost $4000 to clean up, while NuPlant emits 8 tons, each of which would cost $3000 to clean. The town wants to reduce pollution to a maximum of 6 tons. For each of the following, indicate what the total cost of cleanup would be for the firms (note: the firms will only clean up as much as they are required to—they are not running a charity; the town isn’t interested in spending taxpayer money on this, either).

a. The town limits each firm to a maximum of 3 tons of pollution.

b. The town gives each firm 3 tradable pollution permits.
Answer #1

a)

Monolyte will be required to abate 2 (5-3) tons of emission,

Abatement cost for Monolyte=2*4000=$8000

NuPlant will be required to abate 5 (8-3) tons of emission,

Abatement cost for Nuplant=5*3000=$15000

Total Cost of cleanup=8000+15000=$23000

b)

We observe that abatement cost is lower in case of NuPlant.

Monolyte has 3 permits, it will be required to abate 2 tons of emission. It will buy 2 permits from NuPlant at a price beteen $3000 and $4000.

After permits are traded, Nuplant will have 1 pemit left and Nuplant will require to abate 7 tons of emissions.

Abatement cost for Nuplant=Total abatement cost=7*3000=$21000